OpenStudy (anonymous):

What is the simplified form of x^4/y^3/x^5/y^7?

OpenStudy (anonymous):

that should help you out if not ill explain it to you

terenzreignz (terenzreignz):

...... \[\Huge \frac{\color{white}{\text{ha}}\frac{x^4}{y^3}\color{white}{\text{ha}}}{\frac{x^5}{y^7}}\]

OpenStudy (anonymous):

An algebraic expression is a combination of numerals, variables and arithmetic operations such as +, -, * and /. Simplification is one of the major operation of math and to work out problem is very difficult without having clear concept. Simplification is applied to different concepts and these are: simplifying radical expressions, square roots, rational expressions, fractions, exponents, algebraic expressions, complex fractions, equations and expressions. Simplifying Algebraic Expressions with Exponents Exponentiation is a mathematical operations, written as an, involving two numbers, the base a and the exponent n. Let us solve 10x2. 5x3 Given 10x2. 5x3 10x2. 5x3 = (10 * 5) x2 + 3 = 50 x5

OpenStudy (anonymous):

now i think we've given you enough help to solve this so cya